# 2Baidu Inc

### v1

@@ 3.1 Preliminaries
$\mathcal{L}_{LDM}:=\mathbb{E}_{x\sim\sigma(x),y,\epsilon\sim\mathcal{N}}\left[||\epsilon-\epsilon_{\theta}(x_{t},t,\tau(y)||^{2}\right]$ (1)
Where $\epsilon$ is the target noise, $\epsilon_{\theta}(\cdot)$ is a denoising network (i.e., U-Net) to predict noise adding to the latents. $t$ is timestep in diffusion process, $x_{t}$ is the noisy latents in timestep $t$. $\tau(\cdot)$ is the text encoder, where the pretrained CLIP [29] text encoder in Stable Diffusion [33]. During inference, a random Gaussian noise $x_{T}\sim\mathcal{N}(0,1)$ is iterative denoised to $x_{0}$, and the final image is obtained through the decoder $\hat{I}_{0}=\mathcal{D}(x_{0})$.
@@ 3.2 DPG framework for T2I Personalization
Next, we outline the main formulation of our DPG framework.
The deterministic policy applies the action that maximizes the Q-function $Q_{\phi}(\cdot)$, and the Q-function is assumed to be differentiable with respect to the action. In our framework, we regard the latent state, the timestep, and the encoded text condition $\{x_{t},t,\tau(y)\}$ as the input, the predicted noise $z_{t}$ as the action, and the text-to-image generation model denoted by $\epsilon_{\theta}(x_{t},t,\tau(y))$ as the policy. We define the policy function in Eq. 2.
$\hat{z}_{t}$ $=\epsilon_{\theta}(x_{t},t,\tau(y))$ (2)
@@
$\max_{\theta}{\mathbb{E}}\left[Q_{\phi}(x_{t},\epsilon_{\theta}(x_{t},t,\tau(y)))\right]$ (3)
@@
While in the diffusion and denoising process, the policy model $\epsilon_{\theta}(\cdot)$ is optimized to minimize Eq. 1 in one timestep. Therefore, minimizing the reconstruction loss encourages policy model $\epsilon_{\theta}(\cdot)$ to make rewards (i.e., minimizing Eq. 1 equals to maximizing Eq. 3 for diffusion model). In this case, the reward function is obtained by Eq. 5.
$r(x_{t},t,\tau(y))=||\epsilon-\epsilon_{\theta}(x_{t},t,\tau(y))||^{2}$ (5)
Here $Q_{\phi}(\cdot)$ works to predict the one step immediate reward in Eq. 5 for the diffusion model $\epsilon_{\theta}(\cdot)$ given $x_{t}$ to estimate the noise in timestep $t-1$.
@@ 3.3 Learning to "Look Forward"
Since the reward in Eq. 5 represents the one-step direct reward, we aim to leverage the nature of the diffusion process to "look forward" to $\hat{x}_{0,t}$. In the diffusion process, the Gaussian noise $z_{t}$ is added into the initial latent $x_{0}$ at timestep $t$, as shown in Eq. 6.
$x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}z_{t}$ (6)
@@
$\hat{x}_{0,t}=\frac{1}{\sqrt{\bar{\alpha}_{t}}}(x_{t}-\sqrt{1-\bar{\alpha}_{t}}\hat{z}_{t})$ (7)
@@
To achieve the objective of "looking forward" to implicitly guide the focus at different denoising states, the reward function in Eq. 5 can be rewritten in Eq. 8 as follows between $x_{0}$ and $\hat{x}_{0,t}$.
$\begin{split}\mathcal{L}&=||\hat{x}_{0,t}-x_{0}||^{2}\\ &=||\frac{1}{\sqrt{\bar{\alpha}_{t}}}(x_{t}-\sqrt{1-\bar{\alpha}_{t}}\hat{z}_{t})-\frac{1}{\sqrt{\bar{\alpha}_{t}}}(x_{t}-\sqrt{1-\bar{\alpha}_{t}}z_{t})||^{2}\\ &=\frac{1-\bar{\alpha}_{t}}{\bar{\alpha}_{t}}||\hat{z}_{t}-z_{t}||^{2}\end{split}$ (8)
With the optimization objective on $\hat{x}_{0,t}$ in Eq. 8, our DPG framework can learn to look forward from $x_{t}$ to $\hat{x}_{0,t}$ to acquire the implicit guidance at different timestep $t$ to enforce appropriate long-term structural consistency between the generated image and the reference images.
In addition, to make utilization of the prior that the Gaussian noise is added gradually over $t$ steps and denoised step-by-step, we accumulated the reward from $t$ to 00 in Eq. 9 to make consistency with the denoising process rather than calculating one-step direct rewards.
$Q_{\phi}(x_{t},\epsilon_{\theta}(x_{t},t,\tau(y)))=\mathcal{L}(x_{t},\tau(y),t)+\gamma Q_{\phi}(x_{t-1},\epsilon_{\theta}(x_{t-1},t-1,\tau(y)))$ (9)
The Q-function $Q_{\phi}(\cdot)$ in Eq. 9 denotes the accumulated expectation of the history reward. Here, we rewrite the reward function with the timestep discount rate to consider the effect of "Look forward" to $\hat{x}_{0,t}$ in the diffusion process. We denote the discount rate with $\gamma$, then we modify the Eq. 3 as follows, where $B$ denotes the batch size.
$\nabla_{\phi}\frac{1}{B}\sum_{B}Q_{\phi}(x_{t},\epsilon_{\theta}(x_{t},t,\tau(y)))$ (10)
@@ 3.4 Learning Complex Reward
To adopt the DINO similarity as reward, we utilize $\mathcal{D}(\cdot)$ to decode the final generation results $\hat{x}_{0,t}$ to the final generated image $\hat{I}$ by using $\hat{I}=\mathcal{D}(\hat{x_{0,t}})$. Then, we obtain the image embeddings from DINO image encoder $\mathcal{X}(\cdot)$ by $\kappa=\mathcal{X}(I)$. Thus, the reward function $r(\cdot)$ can be formulated as Eq. 11:
$r(x_{t})=1-\hat{\kappa}\cdot\kappa$ (11)
@@
Then, we can inject the supervision of unique visual features of the personalized subjects with the original reconstruction reward, as in Eq. 5. Thus we adopt the combination of both the DINO reward and reconstruction reward to work as the complex reward function, controlled with a weight $\lambda$, where $B$ denotes batch size.
$\nabla_{\phi}\frac{1}{B}\sum_{B}(\lambda Q_{\phi}(x_{t},\epsilon_{\theta}(x_{t},t,\tau(y)))+||\epsilon-\epsilon_{\theta}(x_{t},t,\tau(y))||^{2})$ (12)

### v2

@@ 3.1 Preliminaries
$\mathcal{L}_{LDM}:=\mathbb{E}_{x\sim\sigma(x),y,\epsilon\sim\mathcal{N}}\left[||\epsilon-\epsilon_{\theta}(x_{t},t,\tau(y)||^{2}\right],$ (1)
where $\mathcal{L}_{LDM}$ is a squared error loss, $\epsilon$ is the target noise, $\epsilon_{\theta}(\cdot)$ is a denoising network (i.e., U-Net) to predict noise adding to the latents, $t$ is timestep in diffusion process, $x_{t}$ is the noisy latents in timestep $t$, and $\tau(\cdot)$ is the pretrained CLIP [29] text encoder in Stable Diffusion [33]. During inference, a random Gaussian noise $x_{T}\sim\mathcal{N}(0,1)$ is iterative denoised to $x_{0}$, and the final image is obtained through the decoder $\hat{I}_{0}=\mathcal{D}(x_{0})$.
@@ 3.2 DPG framework for T2I Personalization
Next, we outline the main formulation of our DPG framework. The deterministic policy applies the action that maximizes the Q-function $Q_{\phi}(\cdot)$, and the Q-function is assumed to be differentiable with respect to the action. In our framework, We regard the latent state, the timestep, and the encoded text condition $\{x_{t},t,\tau(y)\}$ as the input, the predicted noise $z_{t}$ as the action, and the text-to-image generation model denoted by $\epsilon_{\theta}(x_{t},t,\tau(y))$ as the policy. We define the policy function as follows,
$\hat{z}_{t}$ $=\epsilon_{\theta}(x_{t},t,\tau(y)).$ (2)
@@
$\max_{\theta}{\mathbb{E}}\left[Q_{\phi}(x_{t},\epsilon_{\theta}(x_{t},t,\tau(y)))\right],$ (3)
@@
While in the diffusion and denoising process, the policy model $\epsilon_{\theta}(\cdot)$ is optimized to minimize Eq. 1 in one timestep. Therefore, minimizing the reconstruction loss encourages policy model $\epsilon_{\theta}(\cdot)$ to make rewards (i.e., minimizing Eq. 1 equals to maximizing Eq. 3 for diffusion model). In this case, the reward function is obtained as follows,
$r(x_{t},t,\tau(y))=||\epsilon-\epsilon_{\theta}(x_{t},t,\tau(y))||^{2}.$ (5)
$Q_{\phi}(\cdot)$ works to predict the one step immediate reward in Eq. 5 for the diffusion model $\epsilon_{\theta}(\cdot)$ given $x_{t}$ to estimate the noise in timestep $t-1$.
@@ 3.3 Learning to "Look Forward"
Since the reward in Eq. 5 represents the one-step direct reward, we aim to leverage the nature of the diffusion process to "look forward" to $\hat{x}_{0,t}$. In the diffusion process, the Gaussian noise $z_{t}$ is added into the initial latent $x_{0}$ at timestep $t$ as follows,
$x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}z_{t}.$ (6)
@@
$\hat{x}_{0,t}=\frac{1}{\sqrt{\bar{\alpha}_{t}}}(x_{t}-\sqrt{1-\bar{\alpha}_{t}}\hat{z}_{t}).$ (7)
@@
To achieve the objective of "looking forward" to implicitly guide the focus at different denoising states, the reward function in Eq. 5 can be rewritten between $x_{0}$ and $\hat{x}_{0,t}$ as follows,
$\begin{split}\mathcal{L}&=||\hat{x}_{0,t}-x_{0}||^{2}\\ &=||\frac{1}{\sqrt{\bar{\alpha}_{t}}}(x_{t}-\sqrt{1-\bar{\alpha}_{t}}\hat{z}_{t})-\frac{1}{\sqrt{\bar{\alpha}_{t}}}(x_{t}-\sqrt{1-\bar{\alpha}_{t}}z_{t})||^{2}\\ &=\frac{1-\bar{\alpha}_{t}}{\bar{\alpha}_{t}}||\hat{z}_{t}-z_{t}||^{2}.\end{split}$ (8)
With the optimization objective on $\hat{x}_{0,t}$ in Eq. 8, our DPG framework can learn to look forward from $x_{t}$ to $\hat{x}_{0,t}$ to acquire the implicit guidance at different timestep $t$ to enforce appropriate long-term structural consistency between the generated image and the reference images. Thus, the reward function in Eq. 8 can be used to update $Q_{\phi}(\cdot)$ at step $8$ in Algorithm 1.
In addition, to make utilization of the prior that the Gaussian noise is added gradually over $t$ steps and denoised step-by-step, we accumulated the reward from $t$ to 00 in Eq. 9 to make consistency with the denoising process rather than calculating one-step direct rewards,
$Q_{\phi}(x_{t},\epsilon_{\theta}(x_{t},t,\tau(y)))=\mathcal{L}(x_{t},\tau(y),t)+\gamma Q_{\phi}(x_{t-1},\epsilon_{\theta}(x_{t-1},t-1,\tau(y))).$ (9)
@@ 3.4 Learning Complex Reward
To adopt the DINO similarity as reward, we utilize $\mathcal{D}(\cdot)$ to decode the final generation results $\hat{x}_{0,t}$ to the final generated image $\hat{I}$ by using $\hat{I}=\mathcal{D}(\hat{x_{0,t}})$. Then, we obtain the image embeddings from DINO image encoder $\mathcal{X}(\cdot)$ by $\kappa=\mathcal{X}(I)$. Thus, the reward function $r(\cdot)$ to be estimated by $Q_{\phi}(\cdot)$ can be formulated as follows,
$r(x_{t})=-(1-\hat{\kappa}\cdot\kappa),$ (10)
@@
Then, we can inject the supervision of unique visual features of the personalized subjects with the original reconstruction reward, as in Eq. 5. Thus we adopt the combination of both the DINO reward and reconstruction reward in Eq. 5 to work as the complex reward function, controlled with a weight $\lambda$, where $B$ denotes batch size. The gradient at step $9$ in Algorithm 1 can be refined as follows,
$\nabla_{\theta}\frac{1}{B}\sum_{B}(\lambda Q_{\phi}(x_{t},\epsilon_{\theta}(x_{t},t,\tau(y)))+(-||\epsilon-\epsilon_{\theta}(x_{t},t,\tau(y))||^{2})).$ (11)